### v1

@@ 5 Main Results
The simplified MSEs of the unbiased case corresponding to the settings in Corollaries 1-3 are given by $O\left(\frac{d}{n^{2}}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right)$, $O\left(\frac{d(n-t)}{tn}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right)$ and $O\left(\frac{d}{(n-c)^{2}}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right)$, respectively.
6 Experiments
We implement CorDP-DME for specific values of $\epsilon,\delta,n,t,c$, and compare the privacy-utility trade-offs against DME with LDP and CDP (with the Gaussian mechanism).***The experiments are performed on commodity hardware/laptop. As the two baselines correspond to unbiased estimates, we compare both biased (see Theorem 1) and unbiased (see Proposition 3) versions of CorDP-DME with LDP and CDP. Fig. 4 shows the privacy-utility trade-offs of different DP-DME cases corresponding to $n=100$, $\delta=10^{-5}$ and $d=20$. Recall that $t$ and $c$ are the thresholds on the minimum responding users and maximum colluding users, respectively. CorDP-DME coincides with CDP when no dropouts or colluding users are considered (left). When dropouts and/or colluding users are considered, the MSE of CorDP-DME lies in between the MSEs of (Gaussian) LDP and CDP (middle/ right). We also show that for many cases (i.e., $t=80$, $c=20$ on the right), CorDP-DME outperforms PrivUnitG [20], which is an approximation of PrivUnit [22], that is proven to be the optimum among all LDP mechanisms that result in unbiased estimates.
@@ 6 Experiments
Figure 4: Privacy-utility trade-offs with no dropouts and no colluding users (left), either dropouts or colluding users (middle), both dropouts and colluding users (right).
7 Conclusions and Limitations
@@
In this work, we present CorDP-DME, a novel differentially private DME protocol that uses correlated Gaussian noise to achieve a favorable balance between utility, resilience to dropouts, and robustness against colluding users. CorDP-DME spans the spectrum between DME with LDP, which provides strong resilience but poor utility, and SecAgg-based approaches, which achieve high utility but have limited resilience to dropouts and collusion. A key insight is that carefully tuned (anti) correlated noise can significantly improve utility compared to independent noise mechanisms, even in adversarial settings with dropouts and collusions. One limitation of CorDP-DME in its current form is that it operates in the continuous domain using Gaussian noise. While this allows for rigorous theoretical analysis, it may present challenges for direct practical implementation such as communication inefficiencies and vulnerabilities to floating point attacks [41, 42]. An important direction for future work is to develop a discrete or quantized variants of CorDP-DME, as well as extending the protocol to non-additive mechanisms.
In terms of theoretical extensions, CorDP-DME opens up several intriguing questions. Is Gaussian noise optimal in this correlated noise setting, or can other noise distributions yield further improvements? How can the noise correlations be optimally designed for different network topologies? Studying these questions may yield further insights and advances.
Overall, CorDP-DME takes an important step towards bridging the gap between LDP and SecAgg-based protocols. The correlated Gaussian noise framework provides a foundation for the future design of DP mechanisms that can achieve high accuracy while maintaining robustness in distributed settings. Our information-theoretic analysis of CorDP-DME will inspire further correlated noise mechanisms for distributed privacy-preserving learning.
@@ Appendix A Generalized Privacy Definition
In this section, we show that the privacy definitions used in DME with LDP and SecAgg based distributed DP are special cases of the proposed privacy constraint in Definition 1.
DME with LDP:
@@
SecAgg based DME with distributed DP:
@@
user $i$rightarrow server: $\bar{M}(\mathbf{x}_{i})=M(\mathbf{x}_{i})+\sum_{j<i}\mathbf{\bar{S}}_{j,i}-\sum_{j>i}\mathbf{\bar{S}}_{i,j},\quad\forall i\in[1:n]$ (19)
where $\mathbf{\bar{S}}_{i,j}$ are uniformly distributed random variables from a finite field $\mathbb{F}_{p}$. Considering no dropouts and no colluding users, the privacy constraint in Definition 1 implies,
$\mathbb{P}(\bar{M}(\bm{x}_{i})\in\mathcal{A}|\bar{M}(\mathbf{x}_{j})=y_{j},\forall j\neq i)\leq e^{\epsilon}\mathbb{P}(\bar{M}(\bm{x}_{i}^{\prime})\in\mathcal{A}|\bar{M}(\mathbf{x}_{j})=y_{j},\forall j\neq i)\!+\!\delta,\quad\forall i$ (20)
@@
$\mathbb{P}(\bar{M}(\bm{x}_{i})\in\mathcal{A}|\bar{M}(\mathbf{x}_{j})=y_{j},\forall j\neq i)$
$=\frac{\mathbb{P}(\bar{M}(\bm{x}_{i})\in\mathcal{A},\sum_{j=1}^{n}\bar{M}(\mathbf{x}_{j})\in\mathcal{A}^{\prime},\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})}{\mathbb{P}(\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})}$ (21)
$=\mathbb{P}(\bar{M}(\bm{x}_{i})\in\mathcal{A}|\sum_{j=1}^{n}\bar{M}(\mathbf{x}_{j})\in\mathcal{A}^{\prime},\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})$
@@
$=\frac{\mathbb{P}(\sum_{j=1}^{n}M(\mathbf{x}_{j})\in\mathcal{A}^{\prime})\mathbb{P}(\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})}{\mathbb{P}(\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})}=\mathbb{P}(\sum_{j=1}^{n}M(\mathbf{x}_{j})\in\mathcal{A}^{\prime})$ (24)
where we use $\sum_{i=1}^{n}\bar{M}(\mathbf{x}_{i})=\sum_{i=1}^{n}M(\mathbf{x}_{i})$ and Shannon’s one-time-pad theorem to derive the last two steps. Substituting (24) in (20) (with all $\mathbf{x}_{j}$, $j\neq i$ fixed and $\mathbf{x}_{i}$ on the LHS and $\mathbf{x}_{i}^{\prime}$ on the RHS) gives the privacy constraint in distributed DP with SecAgg.

### v2

@@ IV-A Main Results
The proof of Theorem 2 is given in Appendix E. The simplified MSEs of the unbiased case corresponding to the settings in Corollaries 1-3 are given by $O\left(\frac{d}{n^{2}}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right)$, $O\left(\frac{d(n-t)}{tn}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right)$ and $O\left(\frac{d}{n(n-c)}\frac{\ln(1/\delta)}{\min\{\epsilon,\epsilon^{2}\}}\right)$, respectively.
The results in Theorems 1 and 2 are based on the specific covariance structure employed by the class of distributions in $\mathcal{D}^{G}_{Z}$, i.e., the structure in (15). We will now demonstrate that, for the case of unbiased estimates, this covariance structure is optimal among all possible covariance matrix configurations. It is important to note that the influence of colluding users on the privacy constraint is dependent on the protocol, as the random variables accessible to the server are determined by the protocol’s specific procedures. In Theorem 3, we establish the general converse result, which characterizes the optimal covariance structure in the absence of collusion.
Theorem 3 (Converse - optimal covariance structure)
Let $\tilde{\mathbf{S}}_{\mathcal{U}}=\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}(\mathbf{x}+\tilde{\mathbf{Z}}_{i})$ be an unbiased estimate of $\mathbf{S}_{\mathcal{U}}$, where $\tilde{\mathbf{Z}}_{i}\sim\mathcal{N}(\mathbf{0}_{d},\sigma^{2}_{i}\mathsf{I}_{d})$ for $i\in[1:n]$. Let $[\tilde{Z}_{1,k},\dotsc,\tilde{Z}_{n,k}]^{T}\sim\mathcal{N}(\mathbf{0}_{n},\Sigma)$ for $k\in[1:d]$, where $\tilde{Z}_{i,k}$ is the $k$th coordinate of $\tilde{\mathbf{Z}}_{i}$ and $\Sigma$ is symmetric positive definite. Define the corresponding MSE of $\tilde{\mathbf{S}}_{\mathcal{U}}$ as,
$\mathrm{MSE}(\Sigma)$ $=\max_{\begin{subarray}{c}\mathcal{U}\subseteq\mathcal{U}_{all}\\ t\leq|\mathcal{U}|\leq n\end{subarray}}\quad\sup_{\mathbf{x}_{j_{1}},\dotsc,\mathbf{x}_{j_{|\mathcal{U}|}}\in\mathbb{B}^{d}}$
$\qquad\mathbb{E}\left[\left\|\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}(\mathbf{x}_{i}+\tilde{\mathbf{Z}}_{i})-\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\mathbf{x}_{i}\right\|^{2}\right]$ (28)
$\forall\Sigma=\Sigma^{T},\quad\Sigma\succ 0$ satisfying $(\epsilon,\delta)$-DP in Definition 1 with $\mathcal{U}_{col}=\emptyset$,
$\mathrm{MSE}\left(\frac{1}{n!}\sum_{i=1}^{n!}\Pi_{i}(\Sigma)\right)\leq\mathrm{MSE}(\Sigma)$ (29)
where $\Pi_{i}(\Sigma)=P_{i}\Sigma P_{i}^{T}$ is the $i$-th permutation of $\Sigma$ defined by the permutation matrix $P_{i}$. Moreover, $\tilde{\Sigma}=\frac{1}{n!}\sum_{i=1}^{n!}\Pi_{i}(\Sigma)$ satisfies $(\epsilon,\delta)$-DP in Definition 1 with $\mathcal{U}_{col}=\emptyset$.
The proof of Theorem 3 is given in Appendix G. As $\tilde{\Sigma}$ has all equal diagonal and all equal off diagonal entries, optimizing over the class of distributions $\mathcal{D}^{G}_{Z}$ in (15) yields the optimal MSE among all Gaussian mechanisms (with i.i.d. coordinates that are arbitrarily correlated among users) for unbiased estimates.
Figure 5: Overview of the offline phase of CorDP-DME in a three-user example: Each user $i$ sends secure information to other users $j$ (denoted by $\mathcal{I}_{i,j}$) to determine the common random seeds required for both users $i$ and $j$ to generate the same shared random variable $\mathbf{S}_{i,j}$.
no dropouts, no collusion only collusion only dropouts dropouts and collusion LDP (for comparison)
$t=10$, $c=0$ $t=10$, $c=2$ $t=8$, $c=0$ $t=8$, $c=2$ $t=10$, $c=0$ (best case)
[…]
$\rho_{*}$ $\to-0.111$ $\to-0.111$ $-0.091$ $-0.089$ 00
MSE (biased) $0.166$ $0.199$ $0.554$ $0.598$ $0.665$
MSE (unbiased) $0.199$ $0.248$ $1.242$ $1.488$ $1.988$
TABLE II: Comparison of optimal parameters across varying dropout and collusion thresholds, with $t$ and $c$ denoting the minimum number of responding users, and maximum number of colluding users, respectively.
IV-B Discussion on CorDP-DME
In this section, we illustrate the key ideas used in CorDP-DME through a simple three-user example. With the notation used in Section II, user $i$ sends $M(\mathbf{x}_{i})=\mathbf{x}_{i}+\mathbf{Z}_{i}$ to the server, that computes $\frac{1}{3}\sum_{i=1}^{3}M(\mathbf{x}_{i})=\frac{1}{3}\sum_{i=1}^{3}\mathbf{x}_{i}+\frac{1}{3}\sum_{i=1}^{3}\mathbf{Z}_{i}$ as the estimate of $\frac{1}{3}\sum_{i=1}^{3}\mathbf{x}_{i}$, considering the unbiased case. Assume that we allow for at most one user to drop out in this example. As there are four possibilities for user dropouts, the possible estimates at the server are given by $E_{1}=\frac{1}{2}(\mathbf{x}_{1}+\mathbf{x}_{2}+\mathbf{Z}_{1}+\mathbf{Z}_{2})$, $E_{2}=\frac{1}{2}(\mathbf{x}_{1}+\mathbf{x}_{3}+\mathbf{Z}_{1}+\mathbf{Z}_{3})$, $E_{3}=\frac{1}{2}(\mathbf{x}_{2}+\mathbf{x}_{3}+\mathbf{Z}_{2}+\mathbf{Z}_{3})$, and $E_{4}=\frac{1}{3}(\sum_{i=1}^{3}\mathbf{x}_{i}+\sum_{i=1}^{3}\mathbf{Z}_{i})$ for $\frac{1}{2}(\mathbf{x}_{1}+\mathbf{x}_{2})$, $\frac{1}{2}(\mathbf{x}_{1}+\mathbf{x}_{3})$, $\frac{1}{2}(\mathbf{x}_{2}+\mathbf{x}_{3})$, and $\frac{1}{3}\sum_{i=1}^{3}\mathbf{x}_{i}$, respectively. The corresponding MSEs are given by, $\mathrm{MSE}_{1}=\frac{1}{4}\mathbb{E}[\|\mathbf{Z}_{1}+\mathbf{Z}_{2}\|^{2}]$, $\mathrm{MSE}_{2}=\frac{1}{4}\mathbb{E}[\|\mathbf{Z}_{1}+\mathbf{Z}_{3}\|^{2}]$, $\mathrm{MSE}_{3}=\frac{1}{4}\mathbb{E}[\|\mathbf{Z}_{2}+\mathbf{Z}_{3}\|^{2}]$, and $\mathrm{MSE}_{4}=\frac{1}{9}\mathbb{E}[\|\mathbf{Z}_{1}+\mathbf{Z}_{2}+\mathbf{Z}_{3}\|^{2}]$. Note that for the unbiased case, the MSE does not depend on the $\mathbf{x}_{i}$’s. To minimize the residual error caused by the worst case dropouts, we need to solve:
$\min_{\mathcal{D}_{\mathbf{Z}_{1},\mathbf{Z}_{2},\mathbf{Z}_{3}}}\max\{\mathrm{MSE}_{1},\mathrm{MSE}_{2},\mathrm{MSE}_{3},\mathrm{MSE}_{4}\}$ (30)
under the privacy constraint in (3), where $\mathcal{D}_{\mathbf{Z}_{1},\mathbf{Z}_{2},\mathbf{Z}_{3}}$ represents the joint distribution of $\mathbf{Z}_{1},\mathbf{Z}_{2},\mathbf{Z}_{3}$. Note that this is exactly the main optimization problem in (7) for this example, with $\alpha_{\mathcal{U}}$ fixed at 1 for unbiased estimates. CorDP-DME considers the class of Gaussian distributions with arbitrary correlations among users for $\mathcal{D}_{\mathbf{Z}_{1},\mathbf{Z}_{2},\mathbf{Z}_{3}}$. Let $\mathbf{Z}_{i,j}$ denote the $j$th coordinate of $\mathbf{Z}_{i}$. We show that for each $j$, the optimum covariance matrix of $[\mathbf{Z}_{1,j},\mathbf{Z}_{2,j},\mathbf{Z}_{3,j}]$ has all equal diagonal and all equal off diagonal entries (Theorem 3). Then, minimizing the worst case MSE over $\mathcal{D}_{\mathbf{Z}_{1},\mathbf{Z}_{2},\mathbf{Z}_{3}}$ reduces to minimizing it over $\sigma^{2}$ and $\rho$ as shown in (15). We then show that the maximum MSE in (30) corresponds to any combination of the maximum number of dropouts, i.e., $\mathrm{MSE}_{i}$ for $i=1,2,3$ for this example.
Next, we incorporate the privacy constraint in (3). Note that for this example, assuming no colluding users, the privacy constraint in (3) simplifies to,
$\mathbb{P}(M(\mathbf{x}_{i})\in\mathcal{A}|M(\mathbf{x}_{j}),M(\mathbf{x}_{k}))$
$\quad\leq e^{\epsilon}\mathbb{P}(M(\mathbf{x}_{i}^{\prime})\in\mathcal{A}|M(\mathbf{x}_{j}),M(\mathbf{x}_{k}))+\delta,$ (31)
for all $i,j,k\in\{1,2,3\}$ where $i\neq j\neq k$. By deriving the conditional Gaussian distributions, (IV-B) simplifies to $-2\rho^{2}+\rho\sigma^{2}+\sigma^{2}-2\sigma^{2}_{\epsilon,\delta}\geq 0$ (see Appendix C for the general proof). Analytically solving (30) under this simplified privacy constraint requires finding the optimal $\rho$ for a fixed $\sigma^{2}$, and then analyzing the behavior of (30) with respect to $\sigma^{2}$. (see Appendix C and Theorems 1-2 for details).
The resulting optimum $\sigma^{2}_{*}$ and $\rho_{*}$ in this example define the joint distribution of $\mathbf{Z}_{1},\mathbf{Z}_{2}$, and $\mathbf{Z}_{3}$ that minimizes the residual error in the aggregate under the worst-case dropouts. In this setup, users directly transmit $M(\mathbf{x}_{i})$ to the server, which computes the aggregate in a single step, completing the protocol in one round. Among existing single-round protocols, which are typically LDP-based, CorDP-DME achieves superior accuracy as shown in Section IV-A.
IV-C Example
In this section, we present an example with $n=10$ users with $d=5$, $\epsilon=2$ and $\delta=10^{-5}$, considering different dropout and collusion thresholds. We demonstrate how the optimal parameters from Theorem 1 vary under each scenario. Table II provides a detailed comparison of the optimal parameters and the corresponding minimum MSEs, as obtained from Theorems 1 and 2, across various conditions. Note that in Table II, for the case with no dropouts, the optimal noise variance added to each user’s private vector approaches infinity, regardless of the number of colluding users. This generalizes the two-user scenario discussed in Section III under similar dropout-free conditions. Additionally, the optimal correlation coefficient converges to $-0.111$, the largest magnitude allowed by valid $10\times 10$ covariance matrices of the form (15) with negative correlation. While the optimal parameters for the case of no dropouts with $c=0$ and $c=2$ converge to the same values asymptotically, the resulting MSEs differ because the MSE is a function of $\sigma^{2}(1+\rho(t-1))$, and $\lim_{\sigma^{2}\to\infty}\sigma^{2}(1+\rho(t-1))$ takes different values in each case.
As shown in columns 4 and 5 of Table II, when dropouts are introduced, employing maximum variance and maximum (negative) correlation becomes ineffective, as even a single dropout can lead to substantial residual noise variance, similar to the behavior seen in SecAgg [18, 44]. Nevertheless, negative correlation can still be applied within a single round, resulting in improved performance compared to LDP as seen by the last column (see Section V for a detailed comparison). This is reflected by the fact that the optimal correlation coefficients remain negative (rather than zero) across all cases in Table II (except last column which shows the LDP baseline).
As expected, columns 4 and 5 (with dropouts) show a reduction in noise variance and the magnitude of the correlation coefficient, compared to the case of no dropouts, to mitigate the impact of dropouts on the MSE. Comparing these columns, we observe that for the same dropout threshold, increasing the number of colluding users requires higher noise variance for each user and reduced noise correlation among users to maintain $(\epsilon,\delta)$-DP. This ensures that the server cannot infer too much information about honest users from colluding ones, but at the cost of increasing the MSE, as seen in columns 4-5.
The last column presents the optimal parameters and MSEs for LDP-based DME under the ideal scenario where all users respond and none collude with the server. Even in this best-case scenario, both biased and unbiased estimates perform worse than all cases examined in CorDP-DME (columns 2-5). Notably, LDP employs the lowest noise variance compared to all CorDP-DME cases. Although it may seem counterintuitive to achieve a lower MSE with higher noise variance, this is made feasible through the use of negative correlation, as indicated in the second row.
@@ IV-C Example
Figure 6: Privacy-utility trade-offs with (a) no dropouts and no colluding users, (b) with colluding users and no dropouts, (c) with dropouts and no colluding users (d) with both dropouts and colluding users. $t$: number of users remaining in the system after dropouts, $c$: number of colluding users.
IV-D Noise generation in CorDP-DME
For a given DP-DME setting with $n$ users, $\epsilon,\delta$ privacy parameters and $t,c$ dropout and colluding user thresholds, CorDP-DME first calculates the corresponding optimum noise distribution and decoder parameters from Section IV-A. Each user then generates their specific noise variables, $\mathbf{Z}_{i}$, ensuring that the joint distribution of $(\mathbf{Z}_{1},\dotsc,\mathbf{Z}_{n})$ follows the structure in (15) with the determined optimal parameters. The following steps are followed in the noise generation protocol:
  1. 1.
    Each pair of users $(i,j)$, $i,j\in[1:n]$, $i\neq j$ generates a pairwise random seed using the Diffie-Hellman key exchange, via secure communications through the server.
  2. 2.
    Using the common seed, each pair of users $(i,j)$ samples the same random vector $\mathbf{S}_{i,j}\in\mathbb{R}^{d}$ from $\mathcal{N}(\mathbf{0}_{d},-\rho_{*}\sigma^{2}_{*}\mathsf{I}_{d})$, [^12] We use use negative correlation among the privacy mechanisms of different users. Hence, $-\rho_{*}>0$. where $\mathbf{S}_{i,j}=\mathbf{S}_{j,i}$, and $\mathbf{S}_{i,j}$, $\mathbf{S}_{i^{\prime},j^{\prime}}$ are independent for any $(i,j)\neq(i^{\prime},j^{\prime})$.
  3. 3.
    Each user $i$ independently generates another noise variable $\mathbf{N}_{i}\sim\mathcal{N}(\mathbf{0}_{d},\sigma^{2}_{*}(1+\rho_{*}(n-1))\mathsf{I}_{d})$.
  4. 4.
    The combined noise added to $\mathbf{x}_{i}$, $i\in[1:n]$ is:
    $\mathbf{Z}_{i}=\sum_{j=1,j<i}^{n}\mathbf{S}_{i,j}-\sum_{j=1,j>i}^{n}\mathbf{S}_{i,j}+\mathbf{N}_{i}.$ (32)
    Subsequently, each user $i$, $i\in[1:n]$ sends $M(\mathbf{x}_{i})=\mathbf{x}_{i}+\mathbf{Z}_{i}$ to the server as per (1). Then, the server decodes the mean estimate using (2), based on the optimum decoder calculated using the number of responding users $|\mathcal{U}|$, and the optimum noise parameters $\sigma^{2}_{*},\rho_{*}$.
Fig. 5 illustrates the overview of the noise generation process implemented during the offline phase.
V Experiments
We implement CorDP-DME for specific values of $\epsilon,\delta,n,t,c$, and compare the privacy-utility trade-offs against DME with LDP and CDP (with the Gaussian mechanism). As the two baselines correspond to unbiased estimates, we compare both biased (Theorem 1) and unbiased (Theorem 2) versions of CorDP-DME with LDP and CDP. Fig. 6 shows the privacy-utility trade-offs of different DP-DME cases corresponding to $n=100$, $\delta=10^{-5}$ and $d=20$. Each case was replicated 20 times and the plots show the MSEs with $95\%$ confidence intervals. Recall that $t$ and $c$ are the thresholds on the minimum responding users and maximum colluding users, respectively. CorDP-DME coincides with CDP when no dropouts or colluding users are considered ( Fig. 6(a)). When considering only colluding users (without any dropouts), the MSE of CorDP-DME increases slightly compared to that of CDP (FIg. 6(b)). This is due to the need for higher variance and reduced correlation to ensure $(\epsilon,\delta)$-DP of the honest users. The MSE of CorDP-DME with dropouts (with or without colluding users) lies in between the MSEs of (Gaussian) LDP and CDP (Figs. 6(c) and (d)). We also show that for many cases (all cases considered in Fig. 6), CorDP-DME outperforms PrivUnitG [12], which is an approximation of PrivUnit [11], that is proven to be the optimum among all LDP mechanisms that result in unbiased estimates. However, we point out that PrivUnit is only applicable to the case where all users’ private vectors have a common (fixed) $L_{2}$ norm, while all Gaussian mechanisms including CorDP-DME can accommodate any private vector with a bounded $L_{2}$ norm.
VI Conclusions and Limitations
@@
In this work, we present CorDP-DME, a novel differentially private DME protocol that uses correlated Gaussian noise to achieve a favorable balance between utility, resilience to dropouts, and robustness against colluding users. CorDP-DME spans the spectrum between DME with LDP, which provides strong resilience but poor utility, and SA-based approaches, which achieve high utility but are also high in complexity and overheads. A key insight is that carefully tuned (anti) correlated noise can significantly improve utility compared to independent noise mechanisms, even in adversarial settings with dropouts and collusions. Important directions for future work includes developing discrete or quantized variants of CorDP-DME, exploring sparse covariance structures to characterize the fundamental trade-off between communication complexity and utility (MSE), and extending the concept of noise correlation to DP-DME mechanisms beyond the class of Gaussian mechanisms considered in this work.
@@ Appendix A Generalized Privacy Definition
In this section, we 1) motivate the privacy constraint in Definition 1 by comparing with alternative DP definitions, 2) show that the privacy definitions used in DME with LDP and SA based distributed DP are special cases of the proposed privacy constraint in Definition 1.
Comparison with Alternative DP Definitions
First we restate our privacy definition here for convenience.
Definition 1: Generalized $(\epsilon,\delta)$-DP for DME: Let $\mathcal{D}_{i}=\{\mathbf{x}\}_{j\neq i}\cup\mathbf{x}_{i}$ and $\mathcal{D}^{\prime}_{i}=\{\mathbf{x}\}_{j\neq i}\cup\mathbf{x}_{i}^{\prime}$ be two neighboring datasets that only differ in the vector of user $i$ for any $i\in\mathcal{U}_{all}\setminus\mathcal{U}_{col}$. Let $\mathcal{G}_{i}=\{\{M(\mathbf{x}_{j})\}_{j\in\mathcal{U}_{all}\setminus\{i\}},\{\mathbf{x}_{j},\mathbf{Z}_{j},\mathcal{I}_{j}\}_{j\in\mathcal{U}_{col}}\}$ denote all the random variables observed by the server from all users except user $i$, for any $i\in\mathcal{U}_{all}\setminus\mathcal{U}_{col}$. For a given $\epsilon\geq 0$ and $\delta\in(0,1)$, a DP-DME scheme ensures $(\epsilon,\delta)$-DP if the following is satisfied.
$\mathbb{P}(M(\mathbf{x}_{i})\in\mathcal{A}|\mathcal{G}_{i})\leq e^{\epsilon}\mathbb{P}(M(\mathbf{x}_{i}^{\prime})\in\mathcal{A}|\mathcal{G}_{i})+\delta,$ (33)
$\forall\mathcal{D}_{i},\mathcal{D}^{\prime}_{i}$, $\forall i\in\mathcal{U}_{all}\setminus\mathcal{U}_{col}$ and $\forall\mathcal{A}\subset\mathbb{R}^{d}$ in the Borel $\sigma$-field.
To motivate this privacy definition, consider the following alternative DP definition that follows from the original DP framework. For two neighboring datasets $\mathcal{D}_{i}$ and $\mathcal{D}^{\prime}_{i}$, let $\mathcal{V}_{\mathcal{D}_{i}}=\{M(\mathbf{x}_{i}),\mathcal{G}_{i}\}$ and $\mathcal{V}_{\mathcal{D}^{\prime}_{i}}=\{M(\mathbf{x}^{\prime}_{i}),\mathcal{G}_{i}\}$ denote the random variables that the server observes from the respective datasets. Then, the privacy mechanism $M$ is said to satisfy $(\epsilon,\delta)$-DP if the following is satisfied.
$\mathbb{P}(\mathcal{V}_{\mathcal{D}_{i}}\in\mathcal{Y})\leq e^{\epsilon}\mathbb{P}(\mathcal{V}_{\mathcal{D}^{\prime}_{i}}\in\mathcal{Y})+\delta,$ (34)
$\forall\mathcal{D}_{i},\mathcal{D}^{\prime}_{i}$, $\forall i\in\mathcal{U}_{all}\setminus\mathcal{U}_{col}$ and $\forall\mathcal{Y}\subset\mathcal{J}$, where $\mathcal{J}$ is the domain of $\mathcal{V}_{\mathcal{D}}$. Note that a privacy mechanism $M$ that satisfies the generalized DP constraint in Definition 1 also satisfies (34), which can be directly observed by multiplying both sides of (3) by $\mathbb{P}(\mathcal{G}_{i})$. However, the inverse argument does not hold for all cases, implying that the DP constraint in Definition 1 is stronger than the one in (34).
Definition 1 generalizes the privacy constraints used in DP-DME with LDP [11, 12, 29] and distributed DP with SA [14, 16], and offers a privacy framework to analyze and compare DP-DME mechanisms with arbitrary correlations among users. Next, we prove that LDP and Distributed DP with SA are special cases of Definition 1.
DME with LDP
@@
SecAgg based DME with distributed DP
@@
user $i$rightarrow server: $\bar{M}(\mathbf{x}_{i})=M(\mathbf{x}_{i})+\sum_{j<i}\mathbf{\bar{S}}_{j,i}-\sum_{j>i}\mathbf{\bar{S}}_{i,j},$ (36)
$\forall i\in[1:n]$ where $\mathbf{\bar{S}}_{i,j}$ are uniformly distributed random variables from a finite field $\mathbb{F}_{p}$. Considering no dropouts and no colluding users, the privacy constraint in Definition 1 implies,
$\mathbb{P}(\bar{M}(\boldsymbol{x}_{i})\in\mathcal{A}|\bar{M}(\mathbf{x}_{j})=y_{j},\forall j\neq i)$
$\quad\leq e^{\epsilon}\mathbb{P}(\bar{M}(\boldsymbol{x}_{i}^{\prime})\in\mathcal{A}|\bar{M}(\mathbf{x}_{j})=y_{j},\forall j\neq i)\!+\!\delta,\quad\forall i$ (37)
@@
$\mathbb{P}(\bar{M}(\boldsymbol{x}_{i})\in\mathcal{A}|\bar{M}(\mathbf{x}_{j})=y_{j},\forall j\neq i)$
$=\frac{\mathbb{P}(\bar{M}(\boldsymbol{x}_{i})\in\mathcal{A},\sum_{j=1}^{n}\bar{M}(\mathbf{x}_{j})\in\mathcal{A}^{\prime},\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})}{\mathbb{P}(\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})}$ (38)
$=\mathbb{P}(\bar{M}(\boldsymbol{x}_{i})\in\mathcal{A}|\sum_{j=1}^{n}\bar{M}(\mathbf{x}_{j})\in\mathcal{A}^{\prime},\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})$
@@
$=\frac{\mathbb{P}(\sum_{j=1}^{n}M(\mathbf{x}_{j})\in\mathcal{A}^{\prime})\mathbb{P}(\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})}{\mathbb{P}(\{\bar{M}(\mathbf{(}x)_{j})=y_{j}\}_{j\neq i})}$ (41)
$=\mathbb{P}(\sum_{j=1}^{n}M(\mathbf{x}_{j})\in\mathcal{A}^{\prime})$ (42)
where we use $\sum_{i=1}^{n}\bar{M}(\mathbf{x}_{i})=\sum_{i=1}^{n}M(\mathbf{x}_{i})$ and Shannon’s one-time-pad theorem to derive the last two steps. Substituting (42) in (A) (with all $\mathbf{x}_{j}$, $j\neq i$ fixed and $\mathbf{x}_{i}$ on the LHS and $\mathbf{x}_{i}^{\prime}$ on the RHS) gives the privacy constraint in distributed DP with SecAgg.